OpenStudy (anonymous):

Find the prime factorization of each number. If the number is prime, state this 56. Please show work.

OpenStudy (anonymous):

would be easiest if you recognize \[56=8\times 7\] and \[8=2^3\] so you get \[56=2^3\times 7\]

OpenStudy (saifoo.khan):

\[\Huge \checkmark\]

OpenStudy (anonymous):

otherwise you can make a little tree. maybe saifoo will show it.

OpenStudy (saifoo.khan):



OpenStudy (saifoo.khan):